# Algebraic and symplectic curves of degree 8

### v1

@@ Introduction
Since the eighties the study of the theory of complex analytic and algebraic varieties have been enriched by the study of pseudo-holomorphic and complex symplectic varieties. In the case of curves in the projective plane these new objects are strongly related to braid monodromy, see [Orevkov2000, Kulikov:07, KhKu:03], and can be constructed by suitable deformations of some arrangements of algebraic plane curves.
The starting point of this paper is an unpublished idea of S.Yu. Orevkov, which is explained in [Golla-Starkston:22] and outlined in §2. The main idea is to deform symplectically a tricuspidal quartic (or deltoid) such that the tangent lines to the cusps are not concurrent. The idea of Orevkov, performed in detail by M. Golla and L. Starkston is to apply a standard Cremona transformation in order to obtain an irreducible symplectic curve with a configuration of topological type of singularities which does not exist in the algebraic category.
@@
For $n=2$ the primary goal is to determine all algebraic curves of degree $8$ with $6$ singular points with the topological of $\mathbb{E}_{6}$. Unfortunately, the goal was too ambitious and has not been reached. As it usually happens, the existence of symmetries is helpful and in this paper we determine all such curves fixed by a non-trivial projective automorphism. There is exactly one such curve $C_{8,2}$ (up to projective automorphism, of course) fixed by an involution and four such curves $C_{8,3}^{i}$, $i=1,\dots,r$, invariant by an automorphism of order $3$; there are no more curves invariants by automorphism. These four curves have equations in conjugate number fields $\mathbb{K}_{i}\subset\mathbb{C}$ isomorphic to $\mathbb{Q}[t]/p(t)$ where $p(t)$ is an irreducible polynomial of degree $4$. A main question is if they share topological properties. Two of the roots of $p(t)$ are real and two complex conjugate; in this last case, complex conjugation is a homeomorphism of $\mathbb{P}^{2}$ reversing orientations on the curves. In the general case, most likely these curves are rigid by dimension arguments.
@@
Some proofs need non-straightforward computer algebra steps and rely heavily on computations in Sagemath [Sagemath]. The steps are described in several notebooks located in https://github.com/enriqueartal/SymplecticOctics which can be executed either in a computer with the last version of Sagemath or online using Binder [binder].
@@
This paper is inspirated in the ideas fruitfully discussed in the workshop Complex and symplectic curve configurations organized by M. Golla, P. Pokorav and L. Starkston in Nantes, December 2022, and in particular in [Golla-Starkston:22]. I would like to thank the organizers, the attendants, and specially the team work on braid monodromy and symplectic curves.
@@ 1. The deltoid and its tangents at the singular points
The equation of the curve in the right-hand side of Figure 1 is:
$v^{4}+4(1+u)v^{3}+18uv^{2}-27u^{2}=0.$
We are going to use this property to have a topological model of the curve and its tangents.
@@
Using the techniques of [ACCT], applied to the right-hand side of Figure 1, we obtain the following result. The justification of this figure can be found in the notebook ConstructionSymplecticGroup.
@@
This monodromy can be also computed using Sagemath [Sagemath] with the optional package Sirocco [Marco2016], but in this case it can be done directly.
In the presentation of $G_{\Delta L}$ we may omit the generator $\ell_{\infty}$ using (R9) which comes from the situation at infinity. Actually, Zariski-van Kampen method can be thought to happen in the blow-up of the projection point (the point at infinity of the vertical lines), see Figure 2. Then (R9) comes from the boundary of a neighbourhood of the exceptional divisor $E$, see [Mumford1961, ACMat:2020]. Note that the natural meridian $e$ of $E$ is the inverse of $c$. The normal crossing situation implies that $e$ (and hence $c$ and $\ell_{1}\cdot\ell_{2}\cdot\ell_{\infty}$) commute with $\ell_{1},\ell_{2},\ell_{\infty}$. This is a consequence of (R2)-(R7): $\ell_{1},\ell_{2}$ commute with $c$ as their conjugation action comes from braids.
@@
Note that $G_{\Delta L}$ is a semidirect product $\mathbb{F}_{4}\rtimes\mathbb{F}_{2}$ where $c_{1},\dots,c_{4}$ are the generators of normal subgroup $\mathbb{F}_{4}$, $\ell_{1},\ell_{2}$ are the generators of $\mathbb{F}_{4}$ and (R2)-(R7), determine the conjugation action.
This group has been computed in [ADT:10], but we need the above computation both for completeness and to deal with the symplectic deformations.
@@ 2. Symplectic deformations
In the context of symplectic geometry, it is possible to construct a deltoid for which the pseudo-holomorphic tangent lines at the cusps are not concurrent. This was communicated long time ago to the author by S.Yu. Orevkov and was formally written in [Golla-Starkston:22, § 8]. Moreover it can be done as a deformation of the algebraic curve which is an isotopy outside a neighbourhood of the triple point.
@@
Using the classical Seifert-van Kampen theorem, the fundamental group of the complement of this symplectic curve has the same presentation of the algebraic one, adding the relations from the situation in the right-hand side of Figure 3, i.e., $[\ell_{1},\ell_{2}]=[\ell_{1},\ell_{\infty}]=[\ell_{2},\ell_{\infty}]=1$. This technique has been used in [ACM:2020b, AC:1998, CatWaj:05] Actually, the following holds.

### v2

@@ Introduction
Since the eighties the study of the theory of complex analytic and algebraic varieties have been enriched by the study of pseudo-holomorphic and complex symplectic varieties. In the case of curves in the projective plane these new objects are strongly related to braid monodromy, see [Ore00, Kul07, KK03], and can be constructed by local deformations of arrangements of algebraic plane curves which can be expressed in terms of braid monodromy factorizations which are locally algebraic.
The starting point of this paper is an unpublished idea of S.Yu. Orevkov, which is explained in [GS22] and outlined in §2. The main idea is to deform symplectically a tricuspidal quartic (or deltoid) such that the tangent lines to the cusps are not concurrent. The idea of Orevkov, performed in detail by M. Golla and L. Starkston is to apply a standard Cremona transformation in order to obtain an irreducible symplectic curve with a configuration of topological type of singularities which does not exist in the algebraic category.
@@
For $n=2$ the primary goal is to determine all algebraic curves of degree $8$ with $6$ singular points with the topological type of $\mathbb{E}_{6}$. Unfortunately, the goal was too ambitious and has not been reached. As it usually happens, the existence of symmetries is helpful and in this paper we determine all such curves fixed by a non-trivial projective automorphism. There is exactly one such curve $C_{8,2}$ (up to projective automorphism, of course) fixed by an involution and four such curves $C_{8,3}^{i}$, $i=1,\dots,r$, invariant by an automorphism of order $3$; there are no more curves invariant by automorphism. These four curves have equations in conjugate number fields $\mathbb{K}_{i}\subset\mathbb{C}$ isomorphic to $\mathbb{Q}[t]/p(t)$ where $p(t)$ is an irreducible polynomial of degree $4$. A main question is if they share topological properties. Two of the roots of $p(t)$ are real and two complex conjugate; in this last case, complex conjugation is a homeomorphism of $\mathbb{P}^{2}$ reversing orientations on the curves. In the general case, most likely these curves are rigid by dimension arguments.
@@
Some proofs need non-straightforward computer algebra steps and rely heavily on computations in Sagemath [Ste23]. The steps are described in several notebooks located in https://github.com/enriqueartal/SymplecticOctics which can be executed either in a computer with the last version of Sagemath or online using Binder [PJe18].
@@
This paper is inspirated in the ideas fruitfully discussed in the workshop Complex and symplectic curve configurations organized by M. Golla, P. Pokorav and L. Starkston in Nantes, December 2022, and in particular in [GS22]. I would like to thank the organizers, the attendants, and specially the team work on braid monodromy and symplectic curves.
@@ 1. The deltoid and its tangents at the singular points
The equation of the curve in the right-hand side of Figure 1 is
$v^{4}+4(1+u)v^{3}+18uv^{2}-27u^{2}=0;$
the line at infinity is the tangent line to one of the cusps; the other cusps are $(0,0),(1,-3)$ and they have vertical tangent lines. The vertical lines $u=a$, $a\in\mathbb{R}$, intersect the real part of the curve at two real points (solid curves in the right-hand side of Figure 1) and at two other points $(a,v_{0}(a)\pm\sqrt{-1}v_{1}(a))$, $v_{0}(a)\in\mathbb{R}$, $v_{1}(a)\in\mathbb{R}_{>0}$; the dotted curve in the right-hand side of Figure 1 represents $(a,v_{0}(a))$. This picture provides a topological model of the curve and its tangents.
@@
Using the techniques of [ACCT01], applied to the right-hand side of Figure 1, we obtain the following result. The justification of this figure can be found in the notebook ConstructionSymplecticGroup.
$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c$$(u_{0},v_{0})$$u=u_{0}$$\ell_{1}$$\ell_{2}$$\ell_{\infty}$$(u_{0},v_{0})$$v=v_{0}$
Figure 2. Generators of the presentation in Proposition 1.2. The base point is $(u_{0},v_{0})$, where $10<u_{0}<$ 1 (the $u$-coordinates) of the affine singular points, and $v_{0}\gg 0$.
@@
This monodromy can be also computed using Sagemath [Ste23] with the optional package Sirocco [MR16], but in this case it can be done directly.
In the presentation of $G_{\Delta L}$ we may omit the generator $\ell_{\infty}$ using (R9) which comes from the situation at infinity. Actually, Zariski-van Kampen method can be thought to happen in the blow-up of the projection point (the point at infinity of the vertical lines), see Figure 3. Then (R9) comes from the boundary of a neighbourhood of the exceptional divisor $E$, see [Mum61, ACM20c]. Note that the natural meridian $e$ of $E$ is the inverse of $c$. The normal crossing situation implies that $e$ (and hence $c$ and $\ell_{1}\cdot\ell_{2}\cdot\ell_{\infty}$) commute with $\ell_{1},\ell_{2},\ell_{\infty}$. This is a consequence of (R2)-(R7): $\ell_{1},\ell_{2}$ commute with $c$ as their conjugation action comes from braids.
@@
Note that $G_{\Delta L}$ is a semidirect product $\mathbb{F}_{4}\rtimes\mathbb{F}_{2}$ where $c_{1},\dots,c_{4}$ are the generators of normal subgroup $\mathbb{F}_{4}$, $\ell_{1},\ell_{2}$ are the generators of $\mathbb{F}_{2}$ and (R2)-(R7), determine the conjugation action.
This group has been computed in [ADT10], but we need the above computation both for completeness and to deal with the symplectic deformations.
@@ 2. Symplectic deformations
In the context of symplectic geometry, it is possible to construct a deltoid for which the pseudo-holomorphic tangent lines at the cusps are not concurrent. This was communicated long time ago to the author by S.Yu. Orevkov and was formally written in [GS22, § 8]. Moreover it can be done as a deformation of the algebraic curve which is an isotopy outside a neighbourhood of the triple point.
@@
Using the classical Seifert-van Kampen theorem, the fundamental group of the complement of this symplectic curve has the same presentation of the algebraic one, adding the relations from the situation in the right-hand side of Figure 4, i.e., $[\ell_{1},\ell_{2}]=[\ell_{1},\ell_{\infty}]=[\ell_{2},\ell_{\infty}]=1$. This technique has been used in [ACM20b, AC98, CW05] Actually, the following holds.